Deep analysis on failing actuators:
- 70% don't show clocking problems
- 30% show slight clocking (coherent between them)
- tried to move pushing with fingers the DSP#264-338 (MIR: 390-556). They don't seems to move inside the hole.
- started single capacitor test on DSP338 (MIR 556) (using accumulators):
- around Tue Aug 23 13:30:56 GMT 2011, after slewing up and down few degrees actuators jumped 2m. Single capacitor verified to NOT touch coil.using accumulators and master diagnostics.
- time hist done showing flat reading to 2m
- global counter all runnings
- changed square wave and readings remain wrongs
- pushing and releasing the corresponding distribution board, the signal respectively appeared and disappeared
- pushing again with slight more pressure, the signal appeared and NOT disappeared with slew tests and other push-pull tests on the distribution board.
- continuing with slew test, after few full range slew (zenith - horizon) the problem bad reading appeared in the same angular zone (close zenith)
- even here the pushing of distibution board was effective.
Start test on all actuators: nothing relevant with slew in the MIR 474, 472, 471, 390, 534.
Found jumping signal on act MIR 450.
- pushing distribution board and slewing again give good and bad signal respectively
- found around 60 deg from horizon a position in which was not possible to have good signal with simple movements
- good signal got only untight the distribution board screws and moving the distribution board.
-checked the actuators referring to the same distribution board: not affected by distribution board movement.
- distribution board was dismounted: Michael found possible crack on a sold in the connector of act MIR 450.
The distribution boards managing actuators 556MIR and 450MIR has been dismounted, cleaned and mounted (slight dirt into the sockets getting capsens boards)
The distribution board managing actuators 472MIR has been released removing two screws at the edge and shaked in Z. Same for distribution 424MIR (3 screws removed).
The slew telemetry data were analyzed and we found/confirmed three main issues:
- the distribution board managing act MIR556 is critical, as well as the three not working actuators in the left side distribution board MIR471, 472, 474 (Fig1 and Fig2 and Fig4)
- the distribution board managing actuators MIR450 is critical VULCANO (effect) (Fig3-4)
- a new actuator heavy affecting the neightbours was identified: MIR317 (DSP183) (Fig5)
The system was switched on in order to verify on the acts MIR 556 and MIR 450 if the intervention fixed the problems.
Negative response on MIR 556: the actuator is still affected by DB movements.
The DB#6 (managing MIR 556) was substituted with the spare one (DB#F2).
Now everything is GOOD on MIR 556 with 30 slew tests.(NO FLAT ATTACCHED)
Negative responses for MIR450.
The actuator 556MIR of DB#F6 corresponds to the 528MIR on DB#F5.
The system was reset @ Tue Aug 23 23:53:07 GMT 2011 for temperature protection. The chiller was found to be on but not active.
System power on again.
Test with capacitor the 528 MIR before substituting the DB: capsens reading 160um (VS 77 std...) but stable at slew. Continue tomorrow with the DB substitution.
- Fig1:Distribution board MIR 556:
- Fig2: Feed Forward residuals:
- Fig3: Distribution board MIR 450:
- Fig4: Distribution board MIR 450 and act MIR471, 472, 474
- Fig5: New bad actuator found:
- 23 Aug 2011